OpenStudy (anonymous):

the formula for the area of a trapezoid is 1/2 (a+b)h. apply the distributive property twice to write this formula in another way

OpenStudy (osanseviero):

Ok, help me help you. I want to make you get the answer. Which is the distributive property?

OpenStudy (anonymous):

1/2.......idk

OpenStudy (osanseviero):

it says this: a*(b + c) = ab + ac

OpenStudy (osanseviero):

Do you follow me?

OpenStudy (anonymous):

theres no (c) i equation. no i do not follow

OpenStudy (osanseviero):

The distributive property is a law of algebra. It is a*(b +c) = ab + ac For example, if you have 3 * ( 2 + 4), you can do 3 * (6) and that is 18 Or you can do: (3*2) + (3*4) = 6 + 12 = 18

OpenStudy (osanseviero):

yes?

OpenStudy (anonymous):

ok